OpenStudy (amistre64):

Whats the integral of ln(x+2)? Im thinking its along the same lines as ln(x)...

OpenStudy (amistre64):

x ln(x+2) - x + C .... maybe?

OpenStudy (anonymous):

let u = x + 2

OpenStudy (anonymous):

The way i am looking at it would end up as (1/x+2)

OpenStudy (amistre64):

zbay, thats derivative I think ;)

myininaya (myininaya):

let u=x+2. so du=dx. therefore we have int(lnu)=u(lnu)-u

OpenStudy (anonymous):

<---- I am a big dummy 8(

OpenStudy (amistre64):

when the sign say push, you pull right ;)

myininaya (myininaya):

=(x+2)ln[x+2]-(x+2)+c

OpenStudy (anonymous):

You're on to me already

OpenStudy (amistre64):

myin, thanx; thats what I was thinking...

myininaya (myininaya):

you're welcome

OpenStudy (amistre64):

hows this look?

myininaya (myininaya):

looks fine. I would have just a u substitution. let u=x+2 so du=dx Since u=x+2 then x=u-2 which if we multiply by 6 we have 6x=6(u-2). so we have int(6(u-2)/u,u)=6int(1-2/u,u)=6int(1,u)-6int(2/u,u)=6u-12lnu+C =6(x+2)-12(x+2)+C

myininaya (myininaya):

oops thats suppose to read 6(x+2)-12ln(x+2)+C

OpenStudy (amistre64):

.... i still need a looooottt of practice with integrals ;) thanx

myininaya (myininaya):

you're super smart amistre. where do you go to college?

OpenStudy (amistre64):

down here in florida; at the local community college

OpenStudy (amistre64):

finally get to take calc1 and stats during the summer.... if the class dont get canceled that is

myininaya (myininaya):

mathematics is a huge. i still don't know alot of it, and I have a masters

OpenStudy (amistre64):

:) I am working on my masters in math.... i figure at least 4 more years ..

myininaya (myininaya):

I want to get my phd in applied science. I am going to do a research in cryptography with someone in the computer science department who knows less math than me, so it should be interesting since I don't know alot of computer science. I'm mainly interested in the theory of computer science.

OpenStudy (amistre64):

thats sounds fun :) I had started out thinking of getting a computer degree since I write my own programs... but math is my best subject since I was a kid :)

myininaya (myininaya):

me too